A rectanglar plot of ground measeures 18 ft by 15 ft. A path is to be constructed within the plot so that a 108 square foot garden can be planted with the path surrounding it. Find the width of the path.
Answer by checkley77(12844)

18*15=270
(18-2x)(15-2x)=108
270-30x-36x+4x^2-108=0
4x^2-66x+162=0
2(2x^2-33x+81)=0
2(2x-27)(x-3)=0
x-3=0
x=3 ft. wide for the path.
pROOF:
(18-2*3)(15-2*3)=108
(18-6)(15-6)=108
12*9=108
108=108